OpenStudy (anonymous):

how to factorise x^2-3x-4?

OpenStudy (anonymous):

\[x^2-3x-4=0\] Find two numbers that multiply to be -4 and add to be -3, and those are the solutions you need to answer.

OpenStudy (anonymous):

oduct sum rule Product=-4 Sum=1 =x^2+x-4x-4=0 =x(x+1)-4(x+1)=0 =(x-4)(x+1) (x-4) =0 & (x+1)=0 ans x=4 or x=-1

OpenStudy (anonymous):

@AfricanBrucelee cannot be,cuz this question comes from a partial fraction question and whe i used the values you gave me the answer didn't work out

OpenStudy (anonymous):

(x+1)(x-4)=0->x=-1,x=4

OpenStudy (anonymous):

@Syderitic That seems legit. thanks man!

OpenStudy (anonymous):

good then i'll attempt using another method.......